# Finite presentability of outer automorphism groups of large-type Artin groups

### v1

@@ 4.1 Co-finiteness
The easy case is that $L_{i}^{\prime}$ has $v$ or $u$ as a leaf: in this case we can make the subgraph of $\Gamma^{\prime}$ spanned by $S_{0}^{\prime}\cup S_{i}^{\prime}$ isomorphic to the corresponding subgraph of $\Gamma$ spanned by $S_{0}\cup S_{i}$, and make $\psi_{i}$ the induced isomorphism $s^{\prime}\mapsto s$. We make $K_{i}^{\prime}=L_{i}^{\prime}\cup e$.
Suppose instead that the leaf of $L_{i}^{\prime}$ contained in the orbit of $e$ is $hu$, where $h\in G_{v}$. Write $f$ for the edge of $L_{i}^{\prime}$ with $hu$ as a leaf. Since $\varphi$ is equivariant and $f$ is not collapsed, $G_{f}=G_{\varphi(f)}$ and so is a standard cyclic or spherical dihedral parabolic, since $\varphi(f)$ is an edge of $K$. We will define $K_{i}^{\prime}=h^{-1}L_{i}^{\prime}\cup e$.
@@
Henceforth assume $G_{f}$ is cyclic, generated by $s$, and hence $h^{-1}G_{f}h$ is cyclic generated by $t$, say. Then by Lemma 2.11, $h^{-1}\in Z_{A_{\Gamma}}(A_{s})\Delta_{s_{0},s_{1}}\Delta_{s_{1},s_{2}}\dots\Delta_{s_{n-1},s_{n}}$, where for each $i$ $s_{i},s_{i+1}$ span an odd parabolic subgroup. Up to left multiplying $h$ by an element of $Z(s)$, we can and henceforth do assume that $h^{-1}=\Delta_{s_{0},s_{1}}\Delta_{s_{1},s_{2}}\dots\Delta_{s_{n-1},s_{n}}$ (this doesn’t disrupt our assumption that $h^{-1}L_{i}$ has $u$ as a leaf, since $Z(s)$ fixes $u$ since $su=u$, or that $h^{-1}th=s$). Note also that $s$ and $t$ are both in $G_{u}\leq G_{v}=A_{S_{0}}$.
We are ready to add edges between the vertices in $S_{i}^{\prime}$ and those in $S_{i}^{\prime}\cup S_{0}^{\prime}$ to build $\Gamma^{\prime}$. If $x^{\prime},y^{\prime}\in S_{i}^{\prime}$, arrange that $m_{x^{\prime}y^{\prime}}=m_{xy}$ (so $A_{S_{i}^{\prime}}$ is isomorphic to $A_{S_{i}}$). Note that all edges from $S_{i}$ to $S_{0}$ in $\Gamma$, by the previous paragraph, have $s$ as their endpoint in $S_{0}$. Take $x\in S_{i}$ and $r\in S_{0}\setminus\{s\}$, and note that they fix points in $T$ on opposite sides of $f$, and $G_{f}=\langle s\rangle$ so neither fixes $f$. A ping pong argument shows that they generate a free group. In $\Gamma^{\prime}$ we slide these edges, that is we arrange that $m_{x^{\prime}t^{\prime}}=m_{xs}$ for each $x^{\prime}\in S_{i}^{\prime}$, and these are all of the edges from $S_{i}^{\prime}$ to $S_{0}^{\prime}$. We define $\psi_{i}$ by $\psi_{i}:x^{\prime}\mapsto h^{-1}xh$ for all $x\in S_{i}$, and as usual $\psi_{i}:r^{\prime}\mapsto r$ for all $r\in S_{0}$. By the structure of $h^{-1}$ as a product of odd dihedral Garside elements, this is a (composition of) the well known edge twist isomorphisms described for instance in [Cri05, Theorem 1], so is in fact an isomorphism.
@@
That $K^{\prime}$ is finite is clear (it has exactly one more edge than $K$). When defining $K^{\prime}$ we showed that it meets no orbit in $T^{\prime}$ more than once, so to show it is a fundamental domain we just have to show it meets every orbit. To see this note that $x\in T^{\prime}$ is any point, then there are two cases. The first is that $\varphi(x)$ is a translate of $\varphi(e)$, which means $x$ is on a translate of $e$ so its orbit intersects $K^{\prime}$. In this other case, $\varphi(x)=gk$ for some $k\in K\setminus\{\varphi(e)\}$, but since $\varphi$ is injective away from the orbit of $e$, the preimage of $k$ is some $k^{\prime}$, which is either in $K^{\prime}$, or its $h^{-1}$ translate is (if we were in the case that $L_{i}$ had $hu$ as an endpoint). Either way the orbit of $k^{\prime}$ intersects $K$, and by equivariance $\varphi(x)=g\varphi(k^{\prime})$ implies $\varphi(x)=\varphi(gk^{\prime})$, and injectivity on these points gives $x=gk^{\prime}$.
In both cases, we have constructed $K^{\prime}$ to be a fundamental domain. If $x\in K^{\prime}$ is a point then $\varphi(x)$ is in the orbit of a point in $K$, but this means $x$ is in the orbit of a point in the preimage of $K$ under $\varphi$ by equivariance. We constructed $K^{\prime}$ to intersect each orbit in $\varphi^{-1}(K)$. Also by construction, $\varphi(K^{\prime})$ meets each orbit in $T$ at most once (we built it out of translates of the $L^{\prime}_{i}$ and all the points with images meeting $A\varphi(e)$, the latter of which were in the preimage of $e$). Since every point in $T$ not in the orbit of $e$ has unique preimage, the only concern is that two points in $e$ are in the same orbit in $T^{\prime}$, but this doesn’t happen since if $gx=y$ for $x,y\in e$ then $g\in G_{\varphi(e)}$ by equivariance, meaning $gv=v$. Any element that setwise fixes $e$ and fixes $v$ fixes $e$ pointwise, so in fact $x=y$.
It is clear that $K^{\prime}$ is finite (it is exactly 1 edge larger than $K$).
Let $\Lambda^{\prime}$ be a chunk of $\Gamma^{\prime}$. We note that, just as $T$ induced a visual splitting of $A_{\Gamma}$, $T^{\prime}$ induces a visual splitting of $A_{\Gamma^{\prime}}$. As such there must exist $i\in\{0\dots n\}$ such that $V(\Lambda^{\prime})\subseteq S_{i}^{\prime}$, as if this were not possible the visual splitting of $A_{\Gamma^{\prime}}$ would induce a visual splitting of $A_{\Lambda^{\prime}}$ over a separating vertex or separating edge, since all edge stabilisers in $T^{\prime}$ are of this form. Now since the induced subgraphs spanned by $S_{i}$ and $S_{i}^{\prime}$ are by definition of $S_{i}^{\prime}$ isomorphic, and by the definition of $\psi$, $\psi(A_{\Lambda^{\prime}})$ is conjugate to a chunk parabolic in $A_{\Gamma}$, so fixes a unique point in $T$. It follows from the definition of the elementary collapse that $A_{\Lambda^{\prime}}$ fixes a in $T^{\prime}$, and this point is unique unless the point fixed by $A_{\Lambda^{\prime}}$ does not have unique preimage under $\varphi$.
@@
4.2 Finitely Presented Stabilisers
In this subsection, we show that the stabilisers of simplices in $\mathcal{F}$ are finitely presented. We show this making use about these stabilisers from [Lev05].
We will first restrict our attention to vertices. Given an $A$-tree $(T,\Omega)$, which we shall as usual often write $T$, we denote the stabiliser of the corresponding point in $\mathcal{F}$ by $\mathrm{Out}^{T}(A)$. If $[\varphi]\in\mathrm{Out}^{T}(A)$, it means by definition there is an equivariant isometry $f:(T,\Omega)\rightarrow(T,\Omega\circ\varphi)$. We write $\mathrm{Out}^{T}_{0}(A)\leq\mathrm{Out}^{T}(A)$ for the subgroup of $\mathrm{Out}^{T}(A)$ where this isometry is moreover $A$-orbit-preserving (note that the $A$-orbit equalivalence relation for $\Omega$ and $\Omega\circ\varphi$ is the same). Observe that $\mathrm{Out}^{T}_{0}(A)$ is the kernel of the induced action of $\mathrm{Out}^{T}(A)\curvearrowright T/A$, (this is well defined because of course inner automorphisms, which act like elements of $A$, act in an $A$-orbit preserving way) and therefore $\mathrm{Out}^{T}_{0}(A)\leq_{f.i.}\mathrm{Out}^{T}(A)$. It is, therefore, enough to show that $\mathrm{Out}^{T}_{0}(A)$ is finitely presented.
We write $(V,E)$ for the graph underlying the graph of groups $T//A$, and $G_{s}$ for the associated simplex groups where $s\in V\cup E$. Following [Lev05] we consider the homomorphism
$\rho:\mathrm{Out}^{T}_{0}(A)\rightarrow\Pi_{v\in V}\mathrm{Out}(G_{v}),$
where on each factor of the co-domain, $\rho(\varphi)$ corresponds to the automorphism $[\varphi]$ induces on the corresponding vertex groups. This definition makes sense because, by our restriction to the subgroup $\mathrm{Out}^{T}_{0}(A)$, there is an element of $[\varphi]$ fixing each vertex of $T$ (and thus inducing an automorphism on $G_{v}$), and this element is well defined up to an element of $\mathrm{Inn}(G_{v})$, so is a well defined class in $\mathrm{Out}(G_{v})$.
We will now show that the $\mathrm{Out}^{T}_{0}(A)$ is finitely presented as the kernel and image of $\rho$ are. We will use the structure of $T$ as a $\Gamma$-tree provided by Lemma 4.10.
@@ 4.2 Finitely Presented Stabilisers
Let $\mathrm{Out}^{T}(A)$ be a point stabiliser in $\mathcal{F}$. Then $\rho(\mathrm{Out}^{T}_{0}(A))$ is finitely presented.
@@
It follows from Lemma 4.10 and Definition 4.3 that every vertex group in $T//A$ is isomorphic to $\mathbb{Z}$, a spherical dihedral Artin group, or a chunk parabolic subgroup. It follows from the fact that $\mathrm{Out}(\mathbb{Z})$ is order 2, Lemma 2.8, and the fact that non-dihedral chunks have finite outer automorphism group (see Definition 3.3) that $\Pi_{v\in V}\mathrm{Out}(G_{v})$ has a finite index subgroup which is abelian and finitely presented. In fact, we may take it to be $\mathbb{Z}^{n}$ where $n$ is the number of vertices of $T//A$ with an even dihedral Artin group as the vertex group. Now any subgroup of $\Pi_{v\in V}\mathrm{Out}(G_{v})$, and in particular $\rho(\mathrm{Out}^{T}_{0}(A))$, is virtually abelian and finitely presented, and so in particular is finitely presented. ∎

### v2

@@ 4.1 Co-finiteness
The easy case is that $L_{i}^{\prime}$ has $v$ or $u$ as a leaf: in this case we can make the subgraph of $\Gamma^{\prime}$ spanned by $S_{0}^{\prime}\cup S_{i}^{\prime}$ isomorphic to the corresponding subgraph of $\Gamma$ spanned by $S_{0}\cup S_{i}$, and make $\psi_{i}$ the induced isomorphism $s^{\prime}\mapsto s$. We make $K_{i}^{\prime}=L_{i}^{\prime}\cup e$. The unlabelled grey edges in Figure 2 represent these cases.
Figure 2: On the left is part of $T$, with solid edges in $K$. On the right is the pre-image under $\varphi$ in $T^{\prime}$.
$\varphi(e)$$v$$u$$e$$he$$K_{i}$$h^{-1}K_{i}$$hu$$L_{i}^{\prime}$$h^{-1}L_{i}^{\prime}$$\varphi$
Suppose instead that the leaf of $L_{i}^{\prime}$ contained in the orbit of $e$ is $hu$, where $h\in G_{v}$. This case is represented by the black edges in Figure 2. Since $he$ is not in $K^{\prime}$, it is clear if we want $K^{\prime}$ to be connected we must add $h^{-1}L_{i}^{\prime}$, instead of $L_{i}^{\prime}$, to $K^{\prime}$.
Write $f$ for the edge of $L_{i}^{\prime}$ with $hu$ as a leaf. Since $\varphi$ is equivariant and $f$ is not collapsed, $G_{f}=G_{\varphi(f)}$ and so is a standard cyclic or spherical dihedral parabolic, since $\varphi(f)$ is an edge of $K$. We will define $K_{i}^{\prime}=h^{-1}L_{i}^{\prime}\cup e$.
@@
Henceforth assume $G_{f}$ is cyclic, generated by $s$, and hence $h^{-1}G_{f}h$ is cyclic generated by $t$, say. Then by Lemma 2.12, $h^{-1}\in Z_{A_{\Gamma}}(A_{t})\Delta_{s_{0},s_{1}}\Delta_{s_{1},s_{2}}\dots\Delta_{s_{n-1},s_{n}}$, where $s_{0}=t$, $s_{n}=s$ and for each $i$ $s_{i},s_{i+1}$ span an odd parabolic subgroup. Note also that $s,t\in G_{v}=A_{S_{0}}$.
We are ready to add edges between the vertices in $S_{i}^{\prime}$ and those in $S_{i}^{\prime}\cup S_{0}^{\prime}$ to build $\Gamma^{\prime}$. If $x^{\prime},y^{\prime}\in S_{i}^{\prime}$, arrange that $m_{x^{\prime}y^{\prime}}=m_{xy}$ (so $A_{S_{i}^{\prime}}$ is isomorphic to $A_{S_{i}}$).
Note that all edges from $S_{i}$ to $S_{0}$ in $\Gamma$ have $s$ as their endpoint in $S_{0}$. To see this take $x\in S_{i}$ and $r\in S_{0}\setminus\{s\}$, and note that they fix points in $T$ on opposite sides of $f$, and $G_{f}=\langle s\rangle$ so neither fixes $f$. A ping pong argument shows that they generate a free group.
In $\Gamma^{\prime}$ we slide these edges, that is we arrange that $m_{x^{\prime}t^{\prime}}=m_{xs}$ for each $x^{\prime}\in S_{i}^{\prime}$, and these are all of the edges from $S_{i}^{\prime}$ to $S_{0}^{\prime}$. We define $\psi_{i}$ by $\psi_{i}:x^{\prime}\mapsto h^{-1}xh$ for all $x\in S_{i}$, and as usual $\psi_{i}:r^{\prime}\mapsto r$ for all $r\in S_{0}$. By the structure of $h^{-1}$ as a product of odd dihedral Garside elements which form a path from $s$ to $t$ (when read from right to left), this is a (composition of) the well known edge twist and Dehn twist isomorphisms described for instance in [Cri05, Theorem 1], so is in fact an isomorphism.
@@
That $K^{\prime}$ is finite is clear (it has exactly one more edge than $K$). It is connected by definition (in particular, the choice of each $K_{i}^{\prime}$ to have either $u$ or $v$ as a leaf).
We show that $K^{\prime}$ is a fundamental domain. By construction it is clear that $\varphi(K^{\prime})$ meets each orbit in $T$ once. By equivariance this passes to $K^{\prime}$ for the points on which $\varphi$ is injective. The only other points are those on $e$ (since $e$ is collapsed to a point in $K$), but since the action is without inversions each of these points is in a different orbit.
It remains to show that $K^{\prime}$ meets every orbit. To see this note that $x\in T^{\prime}$ is any point, then there are two cases. The first is that $\varphi(x)$ is a translate of $\varphi(e)$, which means $x$ is on a translate of $e$ so its orbit intersects $K^{\prime}$. In the other case, $\varphi(x)=gk$ for some $k\in K\setminus\{\varphi(e)\}$, but since $\varphi$ is injective away from the orbit of $e$, the preimage of $k$ is some $k^{\prime}$, which is either in $K^{\prime}$, or its $h^{-1}$ translate is (if we were in the case that $L_{i}$ had $hu$ as an endpoint). Either way the orbit of $k^{\prime}$ intersects $K^{\prime}$, and by equivariance $\varphi(x)=g\varphi(k^{\prime})$ implies $\varphi(x)=\varphi(gk^{\prime})$, and injectivity on these points gives $x=gk^{\prime}$.
Let $\Lambda^{\prime}$ be a chunk of $\Gamma^{\prime}$. We note that, just as $T$ induced a visual splitting of $A_{\Gamma}$, $T^{\prime}$ induces a visual splitting of $A_{\Gamma^{\prime}}$. As such there must exist $i\in\{0\dots n\}$ such that $V(\Lambda^{\prime})\subseteq S_{i}^{\prime}$, as if this were not possible the visual splitting of $A_{\Gamma^{\prime}}$ would induce a visual splitting of $A_{\Lambda^{\prime}}$ over a separating vertex or separating edge, since all edge stabilisers in $T^{\prime}$ are of this form.
By the definitions of $S_{i}^{\prime}$ and $\psi$, $\psi(A_{\Lambda^{\prime}})=h^{-1}A_{\Lambda}h$, where $\Lambda$ is a chunk of $\Gamma$. Since $T$ is a $\Gamma$-tree, $A_{\Lambda}$ fixes a unique point in $K_{i}$, so $\psi(A_{\Lambda^{\prime}})$ fixes a unique point in $h^{-1}K_{i}$. By definition of $K_{i}^{\prime}$, and equivariance of the collapse, $\varphi$ sends $K_{i}^{\prime}$ to $h^{-1}K_{i}$, so $\psi(A_{\Lambda^{\prime}})$ fixes a point in $K^{\prime}$.
It follows from the definition of the elementary collapse that $A_{\Lambda^{\prime}}$ fixes a unique point in $K^{\prime}$ unless the point fixed by $A_{\Lambda^{\prime}}$ does not have unique preimage under $\varphi$.
@@
4.2 A finite index subgroup
We now show how to pass to a finite index subgroup of $\mathrm{Out}(A)$, which we will go on to prove is of type F.
In what follows we say $H\leq A$ is a chunk parabolic if this is true of the image of $H$ under some (equivalently any) isomorphism $A\cong A_{\Gamma}$. This notion is well defined, by chunk rigidity. Note that the notion of a standard chunk parabolic depends on the isomorphism, so is not well defined for $A$.
Write $\mathcal{C}$ for the set of conjugacy classes of chunk parabolic subgroups. This finite set $\mathcal{C}$ is permuted by $\mathrm{Aut}(A)$ and this action descends to $\mathrm{Out}(A)$. Write $\mathrm{Out}_{0}(A)$ for the kernel of this action. Clearly $\mathrm{Out}_{0}(A)$ is finite index.
@@ 4.2 A finite index subgroup
Let $H\leq A$ be chunk parabolic subgroup. Then there is a well defined homomorphism $\rho_{H}:\mathrm{Out}_{0}(A)\rightarrow\mathrm{Out}(H)$.
@@
Take $\varphi$ an arbitrary representative of $[\varphi]\in\mathrm{Out}_{0}(A)$. Then there is $g\in A$ such that $\varphi(H)=gHg^{-1}$. We defined $\rho_{H}(\varphi)=[(\iota_{g}^{-1}\circ\varphi)|_{H}]$.
First we show this does not depend on the choice of $g$. Suppose $g_{1}Hg_{1}^{-1}=g_{2}Hg_{2}^{-1}$, then $g_{2}^{-1}g_{1}$ normalises $H$. By Lemma 2.10 $g_{2}^{-1}g_{1}\in H$, so $(\iota_{g_{1}}^{-1}\circ\varphi)|_{H}$ and $(\iota_{g_{2}}^{-1}\circ\varphi)|_{H}$ differ by an inner automorphism of $H$.
Now we show the definition did not depend on our choice of $\varphi$. If $\iota_{a}\varphi\in[\varphi]$ (where $a\in A$) then $\iota_{a}\varphi(H)=agH(ag)^{-1}$. Clearly $\iota_{g}^{-1}\varphi$ and $\iota_{ag}^{-1}\iota_{a}\varphi$ are the same automorphism, so in particular restrict to the same automorphism of $H$.
Finally to check that $\rho$ is a homomorphism, observe that if $\varphi(H)=g_{1}Hg_{1}^{-1}$ and $\psi(H)=g_{2}Hg_{2}^{-1}$, then one easily checks that $\psi\varphi(H)=(\psi(g_{1})g_{2})H(\psi(g_{1})g_{2})^{-1}$, and $\iota_{g_{2}}^{-1}\psi\circ\iota_{g_{1}}^{-1}\varphi=\iota_{\psi(g_{1})g_{2}}^{-1}\varphi\psi$. ∎
Given $H$ and a conjugate $gHg^{-1}$, the isomorphism $h\mapsto ghg^{-1}$ induces an isomorphism on the outer automorphism groups. This isomorphism does not depend on the choice of $g$, for suppose that $gHg^{-1}=g^{\prime}Hg^{\prime-1}$. By Lemma 2.10 $g^{\prime}=gh$ where $h\in H$. Then take $\varphi:H\rightarrow H$, and note that the images of $[\varphi]$ under the associated isomorphisms are $[\iota_{g}\varphi\iota_{g}^{-1}]$ and $[\iota_{gh}\varphi\iota_{gh}^{-1}]$. Now it suffices to observe that
$(\iota_{g}\varphi\iota_{g}^{-1})^{-1}\circ\iota_{gh}\varphi\iota_{gh}^{-1}$ $=\iota_{g}\varphi^{-1}\iota_{g^{-1}gh}\varphi\iota_{gh}^{-1}$
$=\iota_{g\varphi^{-1}(h)h^{-1}g^{-1}},$
where $\varphi^{-1}(h)$ is well defined as $h\in H$. This is an inner automorphism of $gHg^{-1}$.
We now show that $\rho_{H}$ actually depends only on the conjugacy class of $H$, where $\mathrm{Out}(H)$ and $\mathrm{Out}(gHg^{-1})$ are identified via the induced isomorphism just discussed.
Lemma 4.13.
Given $H$ and a conjugate $aHa^{-1}$, and $[\varphi]\in\mathrm{Out}_{0}(A)$, the isomorphism $\alpha:\mathrm{Out}(H)\rightarrow\mathrm{Out}(aHa^{-1})$ induced by $\iota_{a}:H\rightarrow aHa^{-1}$ identifies $\rho_{H}([\varphi])$ and $\rho_{aHa^{-1}}([\varphi])$.
Proof.
Suppose that $\varphi(H)=g_{1}Hg_{1}^{-1}$ and $\varphi(aHa^{-1})=g_{2}Hg_{2}^{-1}$. Then we directly calculate,
$\alpha(\rho_{H}([\varphi]))^{-1}\rho_{aHa^{-1}}([\varphi])$ $=[(\iota_{a}\iota_{g_{1}}^{-1}\varphi\iota_{a}^{-1})^{-1}|_{aHa^{-1}}(\iota_{g_{2}}^{-1}\varphi)|_{aHa^{-1}}]$
$=[(\iota_{a\varphi^{-1}(g_{1}a^{-1}g_{2}^{-1})})|_{aHa^{-1}}].$
Now we notice that, since $a\varphi^{-1}(g_{1}a^{-1}g_{2}^{-1})$ conjugates $aHa^{-1}$ to itself, it follows by Lemma 2.10 that $a\varphi^{-1}(g_{1}a^{-1}g_{2}^{-1})\in aHa^{-1}$, and so $\alpha(\rho_{H}([\varphi]))=\rho_{aHa^{-1}}([\varphi])$ in $\mathrm{Out}(aHa^{-1})$ as required. ∎
Recall we write $\mathcal{C}$ for the finite set of conjugacy classes of chunk parabolic subgroups. We are now define
$\rho:\mathrm{Out}_{0}(A)\rightarrow\Pi_{[H]\in\mathcal{C}}\mathrm{Out}(H),$
where $\rho$ is the product of the maps $\rho_{H}$, and by the lemmas above this is a well defined homomorphism independent of the choice of $H$.
Lemma 4.14.
Suppose $A$ is not an even dihedral Artin group. Then the image of $\rho$ in $\Pi_{[H]\in\mathcal{C}}\mathrm{Out}(H)$ is finite.
Proof.
It is enough to show that the image in each factor is finite. By Definition 3.3, $\mathrm{Out}(H)$ is finite unless $H$ is a chunk parabolic subgroup coming from a single edge with an even label. Suppose $H$ is such a chunk parabolic subgroup.
Let $\Gamma$ be a defining graph and $\psi:A\rightarrow A_{\Gamma}$ an isomorphism. Without loss of generality, assume $\psi(H)=A_{\Lambda}$, where $\Lambda$ is by assumption an even edge and a chunk in $\Gamma$. Let $s$ be a standard generator associated to a vertex of $\Lambda$ which is of valence at least 2 (this must exist since $A$ is connected and not just an edge by assumption), and write $\Lambda^{\prime}$ for another chunk containing $s$. Let $t$ be the standard generator associated to the other vertex in $\Lambda$.
By [Vas23b, Theorem F], $A_{s}$ is sent by any automorphism to $A_{r}$ up to conjugation, where $A_{r}$ is a standard generator. For all automorphisms in $\mathrm{Out}_{0}(A)$, however, it must be that $A_{r}$ is conjugated into both $A_{\Lambda}$ and $A_{\Lambda^{\prime}}$ (since these are chunk parabolics so by definition of $\mathrm{Out}_{0}(A)$ fixed up to conjugation). So in fact, by the main theorem in [BP23], $A_{r}$ is conjugated to either $A_{s}$ or $A_{t}$.
Note there is no path in $\Gamma$ from $t$ to $\Lambda^{\prime}$ except through $s$, as if there were there would be a cycle involving the edge $\Lambda$, contradicting the fact that $\Lambda$ is a chunk. So there is no path from $t$ to a vertex in $\Lambda^{\prime}$ through only odd labeled edges (because $\Lambda$ is by assumption even labeled), so by Lemma 2.12 $A_{t}$ is not conjugate to the cyclic subgroup generated by a standard generator in $A_{\Lambda^{\prime}}$. So $A_{r}$ must be conjugate to $A_{s}$.
It follows that in $\rho(\mathrm{Out}_{0}(A))\subseteq\mathrm{Out}(H)$ is contained in the subgroup fixing one of the standard generators up to conjugation (more presicely the preimage in $\mathrm{Out}(H)$ of this subgroup of $\mathrm{Out}(A_{\Lambda})$, under the isomorphism induced by $\psi|_{H}$). By Lemma 2.9, this is a finite subgroup. ∎
Hence $\ker(\rho)$ is finite index in $\mathrm{Out}_{0}(A)$ which in turn is finite index in $\mathrm{Out}(A)$. We will write $\mathrm{Out}_{0,+}(A)$ for $\ker(\rho)$.
4.3 Type F stabilisers
In this subsection, we show that the subgroup $\mathrm{Out}_{0,+}(A)$ from the previous subsection acts on $\mathcal{F}$ with type F simplex stabilisers.
We will first restrict our attention to vertices. Given an $A$-tree $(T,\Omega)$, which we shall as usual often write $T$, we denote the stabiliser of the corresponding point in $\mathcal{F}$ by $\mathrm{Out}^{T}(A)$. If $[\varphi]\in\mathrm{Out}^{T}(A)$, it means by definition there is an equivariant isometry $f:(T,\Omega)\rightarrow(T,\Omega\circ\varphi)$. We write $\mathrm{Out}^{T}_{0}(A)\leq\mathrm{Out}^{T}(A)$ for the subgroup of $\mathrm{Out}^{T}(A)$ where this isometry is moreover $A$-orbit-preserving (note that the $A$-orbit equalivalence relation for $\Omega$ and $\Omega\circ\varphi$ is the same).
We write $(V,E)$ for the graph underlying the graph of groups $T//A$, and $G_{s}$ for the associated simplex groups where $s\in V\cup E$. Following [BJ96, Lev05] we consider the homomorphism,
$\rho_{T}:\mathrm{Out}^{T}_{0}(A)\rightarrow\Pi_{v\in V}\mathrm{Out}(G_{v}),$
where on each factor of the co-domain, $\rho_{T}([\varphi])$ corresponds to the automorphism $[\varphi]$ induces on the corresponding vertex groups. This definition makes sense because, by our restriction to the subgroup $\mathrm{Out}^{T}_{0}(A)$, there is an element of $[\varphi]$ fixing each vertex of $T$ (and thus inducing an automorphism on $G_{v}$), and this element is well defined up to an element of $\mathrm{Inn}(G_{v})$, so is a well defined class in $\mathrm{Out}(G_{v})$.
We write $\mathrm{Out}_{0,+}^{T}(A)$ for $\ker(\rho_{T})$. The following lemma shows that, as the name suggests, this is the point stabiliser of $T\in\mathcal{F}$ under the action of the group $\mathrm{Out}_{0,+}^{T}(A)$ from the previous section.
Lemma 4.15.
Given $T$ an $A$-tree in $\mathcal{F}$, $\mathrm{Out}_{0,+}(A)\cap\mathrm{Out}^{T}(A)=\mathrm{Out}_{0,+}^{T}(A)$.
Proof.
We first notice that $\mathrm{Out}^{T}_{0}(A)=\mathrm{Out}^{T}(A)\cap\mathrm{Out}_{0}(A)$, since for $[\varphi]\in\mathrm{Out}^{T}(A)$, $[\varphi]\in\mathrm{Out}^{T}_{0}(A)$ if and only if and only if it preserves the orbits, but this occurs if and only if $[\varphi]$ preserves the vertex orbits (as $T/A$ is a tree, and in particular simplicial), and this is equivalent to the conjugacy classes of the chunk parabolic subgroups being preserved.
Now it is clear that $\rho|_{\mathrm{Out}^{T}_{0}}=\rho_{T}$. The result follows. ∎